A swinging pendulum has a period of 17.0s when the length is shortened by 1.5m, it's period is 8.55 calculate the original length of the pendulum


1
Expert's answer
2022-02-23T08:31:32-0500

The period of a simple pendulum is given by

"T=2\pi\sqrt{l/g}"

So

"T_0=2\pi\sqrt{l_0/g}=17.0\\
T=2\pi\sqrt{(l_0-1.5)/g}=8.55""\frac{l_0}{l_0-1.5}=3.95"

"l_0=3.95l_0-5.93"

"l_0=2.0\:\rm m"